Hi, The gas-phase reaction of NO with F2 to form NOF and F has an activation energy of Ea = 6.30 kJ/mol and a frequency factor of A = 6.00×108 M−1⋅s−1 . The reaction is believed to be bimolecular: NO(g)+F2(g)→NOF(g)+F(g).What is the rate constant at 685 ∘C ?
1
Expert's answer
2020-06-08T15:41:37-0400

According to the Arrhenius equation, the rate constant of the reaction is:

"k = A·e^{-\frac{E_a}{RT}}" ,

where "A" is the frequency factor (more precisely, it is the pre-exponential factor), "E_a" is the activation energy, "R" is the ideal gas constant (8.314 J/(K·mol)) and "T" is the absolute temperature (in kelvin, 685°C is (685+273.15) = 958 K). Therefore:

"k = 6.00·10^8·e^{-\frac{6.30·10^3}{8.314·958}} = 2.72·10^8" M-1s-1.

Answer: the rate constant is 2.72·108 M-1s-1.